OpenStudy (anonymous):

How many solutions does the equation 4x + 8 = −2 − 2 + 5x have?

OpenStudy (jadeishere):

well, first you have to solve the equation 1) isolate the variable 4x + 8 = -2-2+5x 4x = -8-2-2+5x -x = -8-2-2 then -x = -12 x = 12 so one solution >.<

OpenStudy (anonymous):

thank youuuu (:

OpenStudy (jadeishere):

Yep :) \(\LARGE\color{orange}{Welcome~to~Openstudy!}\)